OpenStudy (anonymous):

Find the circumference of a circle whose area is 100m^2

OpenStudy (anonymous):

A = PI * r^2 100 = pi * r^2 r is around 5.641895835 meters C = 2 * pi * r So the circumference is around: 35.449007702 meters

OpenStudy (anonymous):

C = 2pi R and A = pi r^2, find R then substitute it in C

OpenStudy (anonymous):

off to bed now z_z